When a mixture of 14.0 g of acetylene (C2H2) and 18.5 g of oxygen gas is ignited, the combustion reaction produces carbon dioxide and water. (a) Write a balanced chemical equation for the reaction. (b) Which is the limiting reactant? (c) How many grams of C2H2, O2, CO2, and H2O are present after the reaction is complete?
1
Expert's answer
2016-12-15T07:40:10-0500
1) 2C2H2(g) + 5O2(g) --> 4CO2(g) + 2H2O(g)

2) Number of moles of C2H2 = 14 / 26 = 0.5384 mol

Number of moles of O2 = 18.5 / 32 = 0.5781 mol

2 moles of C2H2 reacts with 5 moles of O2,
for 0.5384 moles of C2H2, we expect 1.346 moles of O2 to react, but there is only 0.5781 moles of O2 provided, so O2 is limiting reagent.

3) So 2 moles of C2H2 reacts with 5 mol O2,
0.5781 moles of O2 will react with 0.2312 moles of C2H2.

Moles of C2H2 left = 0.5384 - 0.2312 = 0.3072 moles
Mass of C2H2 left = 0.3072 moles x 26= 7.9872g

Moles of O2 left = 0 moles
Mass of O2 left = 0 g